Determine the mass of As2S3 produced if 10.0 g AsCl3 and 13.4 g H2S are reacted according to the following reaction:

2AsCl3 + 3H2S à As2S3 + 6HCl
1
Expert's answer
2020-10-16T09:00:18-0400

2AsCl3 + 3H2S → As2S3 + 6HCl

M(AsCl3) = 181.28 g/mol

n = m/M

n(AsCl3) "= \frac{10.0}{181.28} = 0.055\;mol"

M(H2S) = 34.1 g/mol

n(H2S) "= \frac{13.4}{34.1} = 0.382\; mol"

For every 2 moles of AsCl3 we need 3 moles of H2S.

For 0.055 moles of AsCl3 we need only 0.0825 moles of H2S.

AsCl3 is the limiting reactant.

"\frac{0.055}{2} = 0.0275" moles of As2S3 will be produced from 0.055 moles of AsCl3.

"m = n\times M"

M(As2S3) = 246.02 g/mol

m(As2S3) "= 0.0275\times 246.02 = 6.76\;g"

Answer: 6.76 g